OpenStudy (anonymous):

Cube X has a volume of x,Cube Y has a volume of 8x. If Cube Y has an edge length of y, what is the edge length of Cube X? y/2 y^3 2y ³√y Cannot be determined by the given information.

OpenStudy (anonymous):

The volume of a cube is equal to the edge length cubed, so (edge length)^3. So, volume=(edge)^3. The volume of cube Y is 8 times that of cube X. Set up two equations, one for cube y and one for cube x: 8x=(y)^3 and x=(edge)^3. From there, cube root both sides of each equation to give you \[\sqrt[3]{8x}=\sqrt[3]{y}\] and\[\sqrt[3]{x}=\sqrt[3]{edge}\] Solve the equations for y and (edge) respectively

OpenStudy (anonymous):

³√y/2

OpenStudy (anonymous):

\[\sqrt[3]{8x}=\sqrt[3]{y^3}\] \[\sqrt[3]{x}=\sqrt[3]{edge^3}\] Use these equations, not the ones i gave you up there

OpenStudy (anonymous):